Grade 11Algebra

Solubility product of silver bromide=5*10^-13 the quantity of potassium bromide (120g/mol) to be added to one litre of 0.05M solution of silver nitrate to start PPT. of silver bromide

Abir
8 Years agoGrade 11

1 Answer

rajesh
8 Years ago
To start ppt ksp=kip. ksp=s×s. 5*10^-13=0.05×conc of Br. conc of Br= 10^-11Weigh of KBr=120×10^-11 To start ppt ksp=kip. To start ppt ksp=kip. To start ppt ksp=kip.